OpenStudy (anonymous):

In a right triangle the lengths of the legs are 5 ft and 12 ft. Find the length of the hypotenuse.

OpenStudy (lalaly):

use pythagorus hypotenuse=h \[h^2=5^2+12^2\]

OpenStudy (lalaly):

\[h^2=169\]

OpenStudy (saifoo.khan):

Righto.

OpenStudy (lalaly):

\[h=13\]

OpenStudy (saifoo.khan):

taking roots on both sides